OpenStudy (anonymous):

z=f(xy). Suppose f'(1)=1/3. Find dz/dx at (1/3, 3).

OpenStudy (bahrom7893):

ok so again treat y as a constant: z=f(xy) dz/dx = f ' (x*y) * y (chain rule)

OpenStudy (bahrom7893):

now plug in whatever you know to find dz/dx

OpenStudy (anonymous):

So dz/dx=f'(xy)*dy/dx?

OpenStudy (bahrom7893):

Ok let me break this down to something a little easier. Suppose you have f(x) = Sin(x*9)

OpenStudy (bahrom7893):

So z = Sin(x*9) dz/dx would then be Cos(x*9) * d(x*9)/dx = Cos(x*9) * 9

OpenStudy (bahrom7893):

In our case z = f(x*y) dz/dx = { d[ f(x*y) ]/dx } * d [ x*y ] /dx

OpenStudy (bahrom7893):